How long will it take a vehicle traveling 72 kilometers per hour to go 36 kilometers?


1
Expert's answer
2022-02-27T17:12:18-0500

Ehe time t it takes to cover a distance s wiyh a speed v is t=s/t. In our case s=36 km, v=72 km/h, so t=s/t=36/72=0.5 hour or 30 minutes.